OpenStudy (loser66):

Mike scored 6000 points in game. In the next five rounds he was out of form and hence, could make only a total of 90 points, as a result of which his average fell by 2 points. How many rounds did he play in all? Please, help

OpenStudy (anonymous):

Let x be the number of rounds played while scoring 6000 points. Then, the average score of these rounds is 6000/x. The total number of rounds played is then x+5 and the total number of points scored is 6090. The average for all the rounds is therefore 6090/(x+5) and this average is 2 less than the average score of the first bunch of rounds. Make sense?

OpenStudy (anonymous):

In other words,\[\frac{ 6090 }{ x+5 } = \frac{ 6000 }{ x } - 2\]

OpenStudy (loser66):

I got what you meant. Thanks for explanation. I made a very complicated equation. :)

OpenStudy (anonymous):

You're welcome.